write an equation in slope-intercept form that satifies each condition
slope=3/4, y-intercept=8
slope=0.5, passess through (-1,7)

Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b

Line with slope=3/4 and y-intercept=8 then y = (3/4)x + 8
Line with slope=0.5 and passing through (-1,7)
y = .5x + b |using ordered pair (-1,7) to solve for b
7 = -.5 b
7.5 = b
Y = .5x + 7.5